# Efficient and operational quantifier of divisibility in terms of channel discrimination

### v1

@@ VI Conclusion
Differently from other methods found in the literature, our toolbox can be cast as a single SDP that decides whether a given family of CPTP maps is CP-divisible and returns the passage maps (if any) connecting any two consecutive instants of time. Additionally, by weakening the formulation of our program, we can similarly investigate the P-divisibility of any given discrete quantum dynamics. In the cases where the dynamics is not CP-divisible, upon a small modification, we can also adapt our method and seek for the best (in the diamond-norm) CPTP map that connects two consecutive instants of time. Finally, we emphasize that our method is applied to any given family of CPTP maps, and artificial requirements of dimensionality or invertibility do not plague it. As an illustration of our method, we analyzed two paradigmatic quantum dynamics in the literature, the dephasing and collisional models. In the collisional case, our quantifier shows a clear difference between the CPTP and P divisibility. This enables us to identify two different non-Markovian regimes, called weak (P-divisible) and strong non-Markovian (non-P-divisible) Chruściński and Maniscalco (2014). In the dephasing dynamics, we were able to identify non-divisibility for the case where the map does not have an inverse. Moreover, we could identify large regions of CPTP-divisibility and also pinpoint the dynamics parameters leading to stronger non-divisibility.
@@
We draw attention to refs. Chruściński et al. (2018) and Anand and Brun (2019b). In the first contribution, the authors center attention on those cases where the generators giving rise to dynamical maps are not invertible. They are able to provide necessary and sufficient conditions in which a given quantum dynamics is divisible—their definition of divisibility embraces completely positive, positive, and functional linear divisibility. In the second contribution, the authors define a measure of non-Markovianity based on robustness. Although related to our work, there are some differences worth pointing out. In Ref. Chruściński et al. (2018), results are tailored to continuous quantum dynamics, and the relation between what they obtain and those results involving discrete evolutions is not evident and deserves further analysis—a fact that is emphasized by the authors themselves. In Ref.Anand and Brun (2019b), the situation is similar. In order to keep their problem in the convex regime, the authors had to focus on a restricted subset of all Choi matrices; those in which the consecutive instants of time are sufficiently close together. In comparison, our approach works for discrete families of CPTP maps, but our method is efficient and versatile and can be adapted to numerically investigate continuous evolutions, provided that we discretize the dynamics.
@@
To conclude, we discuss possible venues of investigation hinted by our results. In ref. Lautenbacher et al. (2022) the authors have also considered dephasing dynamics but in the context of finding the optimal recovery map. We conjecture that instead of looking for the exact recovery map, one could get inspired by our ideas and seek the map that would best approximate the reverse map according to some meaningful norm. In this new perspective, the problem would amount to solving the SDP with the final map "$\Lambda_{2|0}$" given by the identity, which corresponds to minimizing the distance between $\Lambda^{\prime}\circ\Lambda$ and the identity map over all CPTP maps $\Lambda^{\prime}$, given an input CPTP map $\Lambda$. If $\Lambda$ is reversible, then $\Lambda^{\prime}=\Lambda^{-1}$; otherwise, we would find its best approximation according to some norm. That is, one could not only determine whether such a recovery map exists but also exhibit its best approximation. In turn, by considering the dual version of the associated SDP, it should be possible to construct explicit witnesses for non-divisibility, that could be particularly relevant to be applied in experimental setups. These are two interesting research venues opened by this work and we hope our results might trigger further research in these directions.
@@ Acknowledgements.
Cristhiano Duarte thanks for the hospitality of the Institute for Quantum Studies at Chapman University. CD has been funded by an EPSRC grant. This research was partially supported by the National Research, Development and Innovation Office of Hungary (NKFIH) through the Quantum Information National Laboratory of Hungary and through the grant FK 135220. This research was also supported by the Fetzer Franklin Fund of the John E. Fetzer Memorial Trust and by grant number FQXi-RFP-IPW-1905 from the Foundational Questions Institute and Fetzer Franklin Fund, a donor advised fund of Silicon Valley Community Foundation. RC acknowledges the Serrapilheira Institute (Grant No. Serra-1708-15763), the Simons Foundation (Grant Number 1023171, RC) and the Brazilian National Council for Scientific and Technological Development (CNPq, Grant No.307295/2020-6). R.N. acknowledges support from Quantera project Veriqtas. N.K.B. acknowledges financial support from CNPq Brazil (Universal Grant No. 406499/2021-7) and FAPESP (Grant 2021/06035-0). N.K.B. is part of the Brazilian National Institute for Quantum Information (INCT Grant 465469/2014-0).
@@
The “if” part. Suppose that $\rho$ verifies (a) and (b) above. We must prove that its isomorphic image $\mathcal{N}\circ\mbox{T}_{A}$ is completely positive and trace preserving. Firstly, given $\sigma_{A}\in\mathcal{D}(\mathcal{H}_{A})$ note that:
$\mbox{Tr}_{B}[\mathcal{N}(\sigma_{A}^{T})]$ $=\mbox{Tr}_{B}\mbox{Tr}_{A}[\rho(\sigma_{A}^{T}\otimes\mathds{1}_{B})]=\mbox{Tr}_{A}\mbox{Tr}_{B}[\rho(\sigma_{A}^{T}\otimes\mathds{1}_{B})]$
@@
So that $\mathcal{N}\circ\mbox{T}_{A}$ is trace-preserving. Now, it remains to prove that $\mathcal{N}\circ\mbox{T}_{A}$ is completely positive. Take $\sigma_{A}\in\mathcal{D}(\mathcal{H}_{A})$, then:
$(\mathcal{N}$ $\circ\mbox{T}_{A})(\sigma_{A})=$
$=\mbox{Tr}_{A}\left[\rho(\sigma_{A}^{T}\otimes\mathds{1}_{B})\right]=\mbox{Tr}_{A}\left[\sqrt{\rho^{T_{A}}}(\sigma_{A}\otimes\mathds{1}_{B})\sqrt{\rho^{T_{A}}}\right]$
$=\sum_{a}(\bra{a}\otimes\mathds{1}_{B})\left(\sqrt{\rho^{T_{A}}}(\sigma_{A}\otimes\mathds{1}_{B})\sqrt{\rho^{T_{A}}}\right)(\ket{a}\otimes\mathds{1}_{B})$
$=\sum_{a}\left[(\bra{a}\otimes\mathds{1}_{B})\sqrt{\rho^{T_{A}}}\right](\sigma_{A}\otimes\mathds{1}_{B})\left[\sqrt{\rho^{T_{A}}}(\ket{a}\otimes\mathds{1}_{B})\right]$ (24)
Defining, $K_{a}:=[(\bra{a}\otimes\mathds{1}_{B})\sqrt{\rho^{T_{A}}}]$ for all $a$, we can rewrite $\mathcal{N}(\sigma_{A})$ as
$\mathcal{N}(\sigma_{A})=\sum_{a}K_{a}(\sigma_{A}\otimes\mathds{1}_{B})K_{a}^{\ast},$ (25)
@@
The “only if” part. Assume that $\mathcal{N}\circ\mbox{T}_{A}$ is completely positive and trace-preserving. For this part, we use the inversion expressed in eq. (6). As a matter of fact,
$\mbox{Tr}_{B}(\rho)$ $=\mbox{Tr}_{B}\left[\sum_{i,j}^{d}\ket{i}\bra{j}\otimes\mathcal{N}(\ket{j}\bra{i})\right]$
$=\sum_{i,j}^{d}\ket{i}\bra{j}\otimes\mbox{Tr}\left[\mathcal{N}\circ\mbox{T}_{A}(\ket{i}\bra{j})\right]$ (26)
@@ Appendix B Proof of Proposition 12
$\mathcal{N}_{3|1}^{AC}=\mathcal{N}_{3|2}^{BC}\circ\mathcal{N}_{2|1}^{AB}:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{C})$ (27)
is a completely positive trace-preserving map arising from the composition of other two CPTP maps. For the sake of comprehension, denote $\rho_{i|j}^{XY}$ as $\mathcal{C}(\mathcal{N}_{i|j}^{XY})$. Then,
$\varrho_{3|1}^{AC}\circ\mbox{T}_{A}$ $=\rho_{3|1}^{AC}=\mathcal{C}(\mathcal{N}_{3|1}^{AC})=\mathcal{C}(\mathcal{N}_{3|2}^{BC}\circ\mathcal{N}_{2|1}^{AB})$
$=\left[\mbox{id}\otimes(\mathcal{N}_{3|2}^{BC}\circ\mathcal{N}_{2|1}^{AB})\right]\left(\sum_{i,j}\ket{i}\bra{j}\otimes\ket{j}\bra{i}\right)$
$=(\mbox{id}\otimes\mathcal{N}_{3|2}^{BC})(\mbox{id}\otimes\mathcal{N}_{2|1}^{AB})\left(\sum_{i,j}\ket{i}\bra{j}\otimes\ket{j}\bra{i}\right)$ (28)
$=(\mbox{id}\otimes\mathcal{N}_{3|2}^{BC})(\rho_{2|1}^{AB})$
$=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\rho_{3|2}^{BC})(\rho_{2|1}^{AB}\otimes\mathds{1}_{C})\right],$
$=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\rho_{3|2}^{BC})(\varrho_{2|1}^{AB}\otimes\mathds{1}_{C})^{T_{A}}\right]$
=$=\left\{\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\rho_{3|2}^{BC})(\varrho_{2|1}^{AB}\otimes\mathds{1}_{C})\right]\right\}^{T_{A}}.$ (29)
@@
$\varrho_{3|1}=\mbox{Tr}_{B}\left[(\mathds{1}_{A}\otimes\varrho_{3|2}^{BC})^{T_{B}}(\varrho_{2|1}^{AB}\otimes\mathds{1}_{C})\right].$ (30)
The "only if" bit follows in complete analogy, and we will not write it down here.
@@ Appendix C Operational significance and SDP characterization of the diamond norm
The diamond norm, or completely bounded trace norm, is a measure of the ability to distinguish between two channels when generic states, possibly entangled with auxiliary degrees of freedom, are provided. It is defined over the space of linear superoperators, i.e. maps of the form $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ that act linearly on the operators in $\mathcal{L}(\mathcal{H})$, and can be expressed as
$\|\Phi\|_{\diamond}=\max_{\begin{subarray}{c}X\in\mathcal{L}(\mathcal{H}_{d}\otimes\mathcal{H})\\ \|X\|_{1}\leq 1\end{subarray}}\|I_{d}\otimes\Phi(X)\|_{1},$ (31)
where $\mathcal{H}_{d}$ is a $d$-dimensional Hilbert space, with $d=\mathrm{dim}(\mathcal{H})$, $I_{d}$ is the identity map on $\mathcal{H}_{d}$, and $\|\bullet\|_{1}$ is the trace norm for linear operators, equal to the sum of the singular values of the operator. In fact, it can be shown that any extension of $\mathcal{H}$ to $\mathcal{H}_{d^{\prime}}\otimes\mathcal{H}$ with dimension $d^{\prime}>d$ cannot perform better than the value for $\|\Phi\|_{\diamond}$ in the maximisation above Watrous (2018). Consequently, it is indifferent to include a maximisation over $d^{\prime}$ on the definition above. It can also be shown that the optimum above can be attained with a rank-1 $X$ of the form $|u\rangle\langle v|$ for some normalised vectors $|u\rangle,\,|v\rangle\in\mathcal{H}_{d}\otimes\mathcal{H}$. For Hermiticity-preserving maps $\Phi$, $|v\rangle=|u\rangle$.
With the definition above, it is clear that the diamond norm is relevant for the task of distinguishing two arbitrary channels $\Psi_{0}$ and $\Psi_{1}$: Assume that an experimenter is able to prepare any state $\rho\in\mathcal{D}(\mathcal{H}_{d^{\prime}}\otimes\mathcal{H})$, of which they can use the subsystem on $\mathcal{H}$ to go through a under-characterised channel $\Psi$. It is only known that the channel $\Psi$ may be $\Psi_{0}$ with probability $p$, or $\Psi_{1}$ with probability $1-p$. Both $\Psi_{0}$ and $\Psi_{1}$ are known, as well as the probability $p$. Only which channel is actually in effect at a time is unknown. Let $\Psi^{\prime}_{0}\coloneqq I_{d^{\prime}}\otimes\Psi_{0}$ and $\Psi^{\prime}_{1}\coloneqq I_{d^{\prime}}\otimes\Psi_{1}$, for the input state $\rho$, the task reduces to measuring the output state and distinguishing between $\rho_{0}\coloneqq\Psi^{\prime}_{0}(\rho)$ and $\rho_{1}\coloneqq\Psi^{\prime}_{1}(\rho)$ with maximum probability.
@@
$\|\Phi\|_{\diamond}=\max_{\rho_{0},\,\rho_{1}\in\mathcal{D}(\mathcal{H})}\|\left(\sqrt{\rho_{0}}\otimes\mathbb{1}_{\mathcal{H}^{\prime}}\right)\,\mathcal{C}_{\Phi}\,\left(\sqrt{\rho_{1}}\otimes\mathbb{1}_{\mathcal{H}^{\prime}}\right)\|_{1},$ (33)
where $\mathcal{C}_{\Phi}\in\mathcal{L}(\mathcal{H}\otimes\mathcal{H}^{\prime})$ is the Choi map corresponding to $\Phi$ [Eq. (6)] (without the transposition). The characterisation above admits a formulation in terms of the following semidefinite programming,
@@
[…]
@@
for $Y_{0},\,Y_{1}\in\mathcal{L}(\mathcal{H}\otimes\mathcal{H}^{\prime})$. The scalar variables $u_{0}$ and $u_{1}$ can be understood as $\|\mathrm{Tr}_{\mathcal{H}^{\prime}}[Y_{0}]\|_{\infty}$ and $\|\mathrm{Tr}_{\mathcal{H}^{\prime}}[Y_{1}]\|_{\infty}$, respectively, since the inequality constraint implies also $Y_{0},\,Y_{1}\geq 0$.

### v2

@@ VI Conclusion
Unlike other methods found in the literature, our toolbox can be cast as a single SDP that decides whether a given family of CPTP maps is CP-divisible and returns the passage maps (if any) connecting any two consecutive instants of time. Additionally, by weakening the formulation of our program, we can similarly investigate the P-divisibility of any given discrete quantum dynamics. In the cases where the dynamics is not CP-divisible, upon a small modification, we can also adapt our method and seek for the best (in the diamond-norm) CPTP map that connects two consecutive instants of time. Finally, we emphasize that our method is applied to any given family of CPTP maps, and artificial requirements of dimensionality or invertibility do not plague it. As an illustration of our method, we analyzed two paradigmatic quantum dynamics in the literature, the dephasing and collisional models. In the collisional case, our quantifier shows a clear difference between the CPTP and P (non-)divisibility. This enables us to identify two different non-Markovian regimes, called weak (P-divisible) and strong non-Markovian (non-P-divisible) [60]. In the dephasing dynamics, we were able to identify non-divisibility for the case where the map does not have an inverse. Moreover, we could identify large regions of CPTP-divisibility and also pinpoint the dynamics parameters leading to stronger non-divisibility.
@@
We draw attention to Refs. [61] and [62]. In the first contribution, the authors center attention on those cases where the generators giving rise to dynamical maps are not invertible. They are able to provide necessary and sufficient conditions in which a given quantum dynamics is divisible—their definition of divisibility embraces completely positive, positive, and functional linear divisibility. In the second contribution, the authors define a measure of non-Markovianity based on robustness. Although related to our work, there are some differences worth pointing out. In Ref. [61], results are tailored to continuous quantum dynamics, and the relation between what they obtain and those results involving discrete evolutions is not evident and deserves further analysis—a fact that is emphasized by the authors themselves. In Ref. [62], the situation is similar. In order to keep their problem in the convex regime, the authors had to focus on a restricted subset of all Choi matrices; those in which the consecutive instants of time are sufficiently close together. In comparison, our approach works for discrete families of CPTP maps, but our method is efficient and versatile and can be adapted to numerically investigate continuous evolutions, provided that we discretize the dynamics.
@@
To conclude, we discuss possible venues of investigation hinted by our results. In Ref. [63], the authors have also considered dephasing dynamics but in the context of finding the optimal recovery map. We conjecture that instead of looking for the exact recovery map, one could get inspired by our ideas and seek the map that would best approximate the reverse map according to some meaningful norm. In this new perspective, the problem would amount to solving the SDP with the final map $\Lambda_{2|0}$" given by the identity, which corresponds to minimizing the distance between $\Lambda^{\prime}\circ\Lambda$ and the identity map over all CPTP maps $\Lambda^{\prime}$, given an input CPTP map $\Lambda$. If $\Lambda$ is reversible, then $\Lambda^{\prime}=\Lambda^{-1}$; otherwise, we would find its best approximation according to some norm. That is, one could not only determine whether such a recovery map exists but also exhibit its best approximation. In turn, by considering the dual version of the associated SDP, it should be possible to construct explicit witnesses for non-divisibility, that could be particularly relevant to be applied in experimental setups. These are two interesting research venues opened by this work and we hope our results might trigger further research in these directions.
@@ Acknowledgements.
Cristhiano Duarte thanks for the hospitality of the Institute for Quantum Studies at Chapman University. CD has been funded by an EPSRC grant. This research was partially supported by the National Research, Development and Innovation Office of Hungary (NKFIH) through the Quantum Information National Laboratory of Hungary and through the grant FK 135220. This research was also supported by the Fetzer Franklin Fund of the John E. Fetzer Memorial Trust and by grant number FQXi-RFP-IPW-1905 from the Foundational Questions Institute and Fetzer Franklin Fund, a donor-advised fund of Silicon Valley Community Foundation. RC acknowledges the Serrapilheira Institute (Grant No. Serra-1708-15763), the Simons Foundation (Grant Number 1023171, RC) and the Brazilian National Council for Scientific and Technological Development (CNPq, Grant No.307295/2020-6). R.N. acknowledges support from the Government of Spain (Severo Ochoa CEX2019-000910-S and TRANQI), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program) and support from the Quantera project Veriqtas. N.K.B. acknowledges financial support from CNPq Brazil (Universal Grant No. 406499/2021-7) and FAPESP (Grant 2021/06035-0). N.K.B. is part of the Brazilian National Institute for Quantum Information (INCT Grant 465469/2014-0).
@@
The “if” part. Suppose that $\rho$ verifies (a) and (b) above. We must prove that its isomorphic image $\mathcal{N}\circ\mbox{T}_{A}$ is completely positive and trace-preserving. Firstly, given $\sigma_{A}\in\mathcal{D}(\mathcal{H}_{A})$ note that:
$\mbox{Tr}_{B}[\mathcal{N}_{B|A}(\sigma_{A}^{T})]$ $=\mbox{Tr}_{B}\mbox{Tr}_{A}[\rho_{B|A}(\sigma_{A}^{T}\otimes\mathds{1}_{B})]$
$=\mbox{Tr}_{A}\mbox{Tr}_{B}[\rho_{B|A}(\sigma_{A}^{T}\otimes\mathds{1}_{B})]$
@@
So that $\mathcal{N}_{B|A}\circ\mbox{T}_{A}$ is trace-preserving. Now, it remains to prove that $\mathcal{N}_{B|A}\circ\mbox{T}_{A}$ is completely positive. Take $\sigma_{A}\in\mathcal{D}(\mathcal{H}_{A})$, then:
$(\mathcal{N_{B|A}}$ $\circ\mbox{T}_{A})(\sigma_{A})=$
$=\mbox{Tr}_{A}\left[\rho_{B|A}(\sigma_{A}^{T}\otimes\mathds{1}_{B})\right]=\mbox{Tr}_{A}\left[\sqrt{\rho_{B|A}^{T_{A}}}(\sigma_{A}\otimes\mathds{1}_{B})\sqrt{\rho_{B|A}^{T_{A}}}\right]$
[…]
$\mathcal{N}_{B|A}(\sigma_{A})=\sum_{a}K_{a}(\sigma_{A}\otimes\mathds{1}_{B})K_{a}^{\ast},$ (28)
@@
The “only if” part. Assume that $\mathcal{N}_{B|A}\circ\mbox{T}_{A}$ is completely positive and trace-preserving. For this part, we use the inversion expressed in eq. (6). As a matter of fact,
@@ Appendix B Proof of Proposition 12
$\mathcal{N}_{2|0}=\mathcal{N}_{2|1}\circ\mathcal{N}_{1|0}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ (30)
is a completely positive trace-preserving map arising from the composition of the other two CPTP maps. For the sake of comprehension, denote $\rho_{i|j}$ as $\mathcal{J}(\mathcal{N}_{i|j})$. Then,
$\varrho_{2|0}\circ\mbox{T}_{\mathcal{H}_{0}}$ $=\rho_{2|0}=\mathcal{J}(\mathcal{N}_{2|0})=\mathcal{J}(\mathcal{N}_{2|1}\circ\mathcal{N}_{1|0})$
[…]
$=(\mbox{id}\otimes\mathcal{N}_{2|1})(\rho_{1|0})$
$=\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\rho_{2|1})(\rho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right],$
$=\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\rho_{2|1})(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})^{T_{\mathcal{H}_{0}}}\right]$
$=\left\{\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\rho_{2|1})(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right]\right\}^{T_{\mathcal{H}_{0}}}.$ (32)
@@
$\varrho_{2|0}=\mbox{Tr}_{\mathcal{H}_{1}}\left[(\mathds{1}_{\mathcal{H}_{0}}\otimes\varrho_{2|1})^{T_{\mathcal{H}_{1}}}(\varrho_{1|0}\otimes\mathds{1}_{\mathcal{H}_{2}})\right].$ (33)
The "only if" direction follows in complete analogy, and we will not write it down here.
@@ Appendix C Operational significance and SDP characterization of the diamond norm
The diamond norm, also known in the literature as the completely bounded trace norm, is a measure of the ability to distinguish between two channels when generic states, possibly entangled with auxiliary degrees of freedom, are provided. It is defined over the space of linear superoperators, i.e. maps of the form $\Phi:\mathcal{L}(\mathcal{H}_{A})\rightarrow\mathcal{L}(\mathcal{H}_{B})$ that act linearly on the operators in $\mathcal{L}(\mathcal{H}_{A})$, and can be expressed as
$\|\Phi\|_{\diamond}=\max_{\begin{subarray}{c}X\in\mathcal{L}(\mathcal{H}_{d}\otimes\mathcal{H}_{A})\\ \|X\|_{1}\leq 1\end{subarray}}\|I_{d}\otimes\Phi(X)\|_{1},$ (34)
where $\mathcal{H}_{d}$ is a $d$-dimensional Hilbert space, with $d=\mathrm{dim}(\mathcal{H}_{A})$, $I_{d}$ is the identity map on $\mathcal{H}_{d}$, and $\|\bullet\|_{1}$ is the trace norm for linear operators, equal to the sum of the singular values of the operator. In fact, it can be shown that any extension of $\mathcal{H}_{A}$ to $\mathcal{H}_{d^{\prime}}\otimes\mathcal{H}_{A}$ with dimension $d^{\prime}>d$ cannot perform better than the value for $\|\Phi\|_{\diamond}$ in the maximisation above [44]. Consequently, it is indifferent to include a maximisation over $d^{\prime}$ on the definition above. It can also be shown that the optimum above can be attained with a rank-1 $X$ of the form $|u\rangle\langle v|$ for some normalised vectors $|u\rangle,\,|v\rangle\in\mathcal{H}_{d}\otimes\mathcal{H}_{A}$. For Hermiticity-preserving maps $\Phi$, $|v\rangle=|u\rangle$.
With the definition above, it is clear that the diamond norm is relevant for the task of distinguishing two arbitrary channels $\Psi_{0}$ and $\Psi_{1}$: Assume that an experimenter is able to prepare any state $\rho\in\mathcal{D}(\mathcal{H}_{d^{\prime}}\otimes\mathcal{H}_{A})$, of which they can use the subsystem on $\mathcal{H}_{A}$ to go through a under-characterised channel $\Psi$. It is only known that the channel $\Psi$ may be $\Psi_{0}$ with probability $p$, or $\Psi_{1}$ with probability 1p1-p. Both $\Psi_{0}$ and $\Psi_{1}$ are known, as well as the probability $p$. Only which channel is actually in effect at a time is unknown. Let $\Psi^{\prime}_{0}\coloneqq I_{d^{\prime}}\otimes\Psi_{0}$ and $\Psi^{\prime}_{1}\coloneqq I_{d^{\prime}}\otimes\Psi_{1}$, for the input state $\rho$, the task reduces to measuring the output state and distinguishing between $\rho_{0}\coloneqq\Psi^{\prime}_{0}(\rho)$ and $\rho_{1}\coloneqq\Psi^{\prime}_{1}(\rho)$ with maximum probability.
@@
$\|\Phi\|_{\diamond}=\max_{\rho_{0},\,\rho_{1}\in\mathcal{D}(\mathcal{H}_{A})}\|\left(\sqrt{\rho_{0}}\otimes\mathbb{1}_{\mathcal{H}_{B}}\right)\,\mathcal{C}_{\Phi}\,\left(\sqrt{\rho_{1}}\otimes\mathbb{1}_{\mathcal{H}_{B}}\right)\|_{1},$ (36)
where $\mathcal{C}_{\Phi}\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ is the Choi map corresponding to $\Phi$ [Eq. (6)] (without the transposition). The characterisation above admits a formulation in terms of semidefinite programming as follows:
@@
s.t. $\quad u_{0}\,\mathbb{1}\geq\mathrm{Tr}_{\mathcal{H}_{B}}[Y_{0}]$
$\quad u_{1}\,\mathbb{1}\geq\mathrm{Tr}_{\mathcal{H}_{B}}[Y_{1}]$
@@
for $Y_{0},\,Y_{1}\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$. The scalar variables $u_{0}$ and $u_{1}$ can be understood as $\|\mathrm{Tr}_{\mathcal{H}_{B}}[Y_{0}]\|_{\infty}$ and $\|\mathrm{Tr}_{\mathcal{H}_{B}}[Y_{1}]\|_{\infty}$, respectively. There is no need to introduce inequalities $\|\mathrm{Tr}_{\mathcal{H}_{B}}[Y_{x}]\|_{\infty}\geq-u_{x}\,\mathbb{1},\,(x=0,1)$, since the last matrix inequality constraint implies also $Y_{0},\,Y_{1}\geq 0$, and thus that their eigenvalues are nonnegative.
More explicitly, one possible standard form for a semidefinite program can be characterized by a triple $(\Xi,C,D)$ [44, 45], such that $\Xi$ is a superoperator mapping operators in $\mathcal{L}(\mathcal{V})$ to operators in $\mathcal{L}(\mathcal{W})$, and $C\in\mathcal{L}(\mathcal{V})$, $D\in\mathcal{L}(\mathcal{W})$ are Hermitian operators acting on the respective vector spaces $\mathcal{V}$ and $\mathcal{W}$. Its primal formulation is given by
$\operatorname*{Maximise}_{X\,\in\,\mathcal{L}(\mathcal{V})}$ $\quad\mathrm{Tr}[C\,X]$
s.t. $\quad\Xi(X)=D,$
$\quad X\geq 0,$ (38)
with a respective dual formulation given by
$\operatorname*{Minimise}_{Y\,\in\,\mathcal{L}(\mathcal{W})}$ $\quad\mathrm{Tr}[D\,Y]$
s.t. $\quad\Xi^{\dagger}(Y)\geq C,$
$\quad Y=Y^{\dagger},$ (39)
with $\Xi^{\dagger}$ the dual map to $\Xi$. The exact optimization sense is unimportant, since it can be changed from minimisation to maximisation and vice-versa by changing the sign of the objective function. The exact prescription for $\Xi$, $C$ and $D$ for the diamond norm can be found in [45]. In the next subsection, the more relevant prescription for the problem introduced in Eq. (IV) will be provided.